# The Power of Linear Programming in Sponsored Listings Ranking: Evidence from Field Experiments

### v2

@@ 4.1 Experiment design
    Benchmark (Control) For consumers assigned to this group, the displayed listings will be generated by the score-based algorithm.
During the 19 days of the experiment, each consumer who entered the sponsored listings page would be randomly assigned into one of three groups where the displayed listings are generated by the corresponding method. The mechanism of assignment adopted by Marketplace A is based on unique consumers’ IDs. It means that within the duration of the experiment, each consumer will be guaranteed to receive the treatment from a fixed group, no matter how many times she re-enters the experiment.
4.2 Experiment statistics and randomization test
In what follows, we present a detailed description of the post-experiment statistics and the randomization test for the pre-experiment data. During the 19-day period, we recorded in total 32,058,321 consumers, of which 10,686,419 consumers were assigned to LP90 (Treatment 1), 10,689,158 consumers were assigned to LP95 (Treatment 2), and 10,682,744 were assigned to Benchmark (Control). The visit that each consumer pays to Marketplace A’s marketplace is called impression. During the experiment period, it is likely that each consumer pays multiple visits, generating many impressions. As a result, in total, 329 million impressions were generated. Specifically, 109,843,211 impressions were assigned to 109,868,303 impressions were assigned to LP90 (Treatment 1), LP95 (Treatment 2), and 109,567,246 consumers were assigned to Benchmark (Control). Panel A of Table 3 demonstrates the above information, as well as the aggregated revenue and gross merchandise volume (GMV), collected within the experiment period. GMV is defined as the total amount of sales completed through Marketplace A’s marketplace over a specified period of time, while revenue is a partial amount of commission fee that Marketplace A takes from GMV.
Then we show some post-experiment statistics and p-values that inspect the randomness of the assignment, that is, whether the impressions were truly randomly assigned without any bias or not. As shown in Panel A of Table 3, for the number of consumers and impressions, the minimum pairwise p-value among the three groups is larger than $0.94$, indicating no significant signs of the differences among all three groups. To further strengthen the randomization of the assignment, we run the t-tests on four weeks’ pre-experiment data for the key metrics of consumers who entered our A/B tests later. Panel B of Table 3 presents information on the average expenditure spent, generated revenue, and the number of purchases made per consumer during the pre-experiment period. We observe that the number of consumers and impressions assigned to each group is statistically indistinguishable, as the minimum pairwise p-values of all pre-experiment tests are larger than $0.22$. The results of the randomization tests indicate that any difference in the consumers’ purchasing behavior among these three experiment groups should be attributed to the proposed approach to which they have been assigned to.
Table 3: Summary Statistics of randomization test
LP90 LP95 Benchmark Min pairwise p-value
(Treatment 1) (Treatment 2) (Control)
Panel A: Summary Statistics for experiment data
Number of consumers 10,686,419 10,689,158 10,682,744 >0.99
Number of impressions 109,868,303 109,843,211 109,567,246 >0.94
Panel B: Randomization Check for pre-experiment data (per consumer)
Expenditure(USD) * * * >0.22
Revenue(USD) * * * >0.94
Number of purchases * * * >0.99
  • Notes. * Due to the confidentiality of data sources, consumer-level data is forbidden to be disclosed by Marketplace A. Nevertheless, we are able to demonstrate the pairwise p-value to show the randomness of the experiment assignment.
5 Field experiment: results and analysis
In this section, we present the main results collected from the field experiment. We start with a detailed description of the financial performance of the participating experiment groups. We present three metrics as the assessments of the ranking quality (revenue and relevance) for experiment groups, which are respectively revenue, purchase count (total number of transactions), and GMV (gross merchandise volume). These are also what Marketplace A adopts as the major evaluation metrics. The purchase count measures the relevance score of the sponsored listings, and it is the realization of purchase probability. The revenue and GMV are metrics for evaluating the revenue of the marketplace. In later parts of this section, we pose a detailed analysis of the experiment data, attempting to better explain the results and interactions of these metrics across all groups. Additionally, we identify the trade-off between relevance and revenue shown by the evidence of the field experiment, which to our knowledge, is the first time in this field of research.
5.1 Performance and analyses
5.1.1 Financial performance
We start by presenting and comparing several core metrics involved in the field experiment, which are all normalized by the number of consumers. Notably, we define and present the comparison (relative change and p-values) among RevenuePerConsumer, GMVPerConsumer, NumberPurchasesPerConsumer, and AverageTransactionPrice, attributed to three experiment groups (LP95, LP90, Benchmark). In particular, RevenuePerConsumer is the total revenue generated per consumer during the experiment period; GMVPerConsumer is the total expenditure spent per consumer during the experiment period; NumberPurchasesPerConsumer is the total number of purchases made per consumer during the experiment period; AverageTransactionPrice is the average item sold price per transaction. Table 4 also demonstrates the relative change of two treatment groups (LP95 and LP90) over the control group (Benchmark) and the associated p-values of the pairwise t-tests.
The results revealed in Table 4 show that, compared to Benchmark, both LP95 and LP90 exhibit improvements in all major metrics, some of which are statistically significant. For LP95, RevenuePerConsumer, GMVPerConsumer, and NumberPurchasesPerConsumer are respectively 0.0293, 0.5553, and 0.0162, indicating 0.81%, 2.96%, and 3.22% increases compared to Benchmark. Among them, increases of GMVPerConsumer and NumberPurchasesPerConsumer are significant with p-values less than 0.001. On the other hand, for LP90, RevenuePerConsumer, GMVPerConsumer, and NumberPurchasesPerConsumer are respectively 0.0295, 0.5465, and 0.0159, indicating 1.80%, 1.39%, and 1.55% increases compared to Benchmark. Amid them, GMVPerConsumer and NumberPurchasesPerConsumer of the LP-based algorithm display significantly distinguishable improvement over Benchmark, with p-values less than 0.03.
Table 4: Relative changes for the financial performance of the LP-based algorithm over the score-based approach
LP95 LP90
(Treatment 1) (Treatment 2)
RevenuePerConsumer (USD) 0.81%  (0.32) 1.80%  (<0.03)
GMVPerConsumer (USD) 2.96%  (<0.001) 1.39%  (0.11)
NumberPurchasesPerConsumer 3.22%  (<0.001) 1.55%  (<0.001)
AverageTransactionPrice (USD) -0.23% -0.20%
5.1.2 Average transaction price
Another critical but less direct metric that we investigate is AverageTransactionPrice. Indeed, AverageTransactionPrice reveals some unique information that accesses the ranking quality from another perspective. To promote the overall revenue, an immediate approach that the ranking model could take is to display listings with higher prices, forcing the consumers to spend more expenditure on the same items. Such shortcuts causing price discrimination will lift the revenue in the short term, yet, it will damage the consumers’ experiences in the longer run, thus hurting the long-term revenue (Nishimura,, 1989). Ideally, we would wish for a ranking model that could recommend more relevant but not necessarily pricier items to consumers. Thus, here we examine AverageTransactionPrice of LP95 and LP90, which are actually 0.23% and 0.20% lower than Benchmark, indicating that the revenue increase obtained by the LP-based approach does not originate from the lifted listing prices.
5.1.3 Regression analysis
[…]
Next, we implement ordinary least squares (OLS) regression model with the following specification:
$Y_{t}=\beta_{0}+\beta_{1}Approach_{t}+\epsilon_{t},$ (13)
where $Y_{t}$ is the dependent variable, and $Y_{t}\in\{\textit{RevenuePerImpression}_{t},\textit{PurchaseIndicator}_{t},\\ \textit{LogAveagePrice}_{t},\textit{WeightedPTR}_{t},\textit{FirstPageRankIndicator}_{t}$}; $\beta_{0}$ is the intercept; $\beta_{1}$ is the coefficient; $Approach_{t}$ is the binary variable indicating which method (treatment or control) the current impression has been assigned to; $\epsilon_{t}$ is the error term. We run this regression with respect to LP90 and LP95, respectively. Due to the exceedingly large scale of full data and practical restrictions in terms of data privacy, we here examine the sample data by uniformly randomly selecting 690,611 impressions for each experiment group, which counts 2,071,833 impressions in total. Since we have shown that the field experiments were truly randomly assigned and according to the law of large numbers, we believe that the regression results on the sampled data should be representative as well. Later on, we will also show more evidence for regression robustness.
Table 5 demonstrates the regression results in impression level. Since Benchmark is picked as the baseline in our experiment, the coefficients shown in Table 5 indicate the difference in attributions to the dependent variables between two LP-based algorithms and Benchmark. Both the coefficients in Column (1) and Column (2) are positive, revealing that the impression assigned to the LP-based algorithms will generate more revenue and purchases than Benchmark. Furthermore, Columns (3) and (4) suggest the reason for LP90’s and LP95’s better performances are not due to the lifted price, since the coefficients in Column (3) are negative, indicating that the LP-based methods do not recommend pricier items. Instead, Column (4) indicates that such an increase in performance can be attributed to the fact that LP-based methods display more relevant items. Column (5) also strengthens the results by showing the positive coefficient of LP95 alongside a marginally negative coefficient of LP90. It suggests that, compared to Benchmark, more items were purchased in the top $5$ slots under LP95, since LP95 is expected to produce more relevant ranked listings.
Table 5: OLS regression results on the treatment effect
Dependent varaible
RevenuePerImpression PurchaseIndicator LogAveagePrice WeightedPTR FirstPageRankIndicator
(1) (2) (3) (4) (5)
LP90 0.056**** $0***$064**** 0.1169**** -0.003****
[0.016] [0.000] [0.002] [0.061] [0.001]
LP95 0.120**** 0.014**** -0.1091**** 0.4826**** 0.0116****
[0.017] [0.000] [0.002] [0.074] [0.001]
Observations 2,071,833 2,071,833 2,071,833 2,071,833 2,071,833
  • Notes. The value inside the square brackets is the robust standard error.
    **** $p<0.001$.
5.2 Heterogeneous treatment effects analysis
In this section, we present heterogeneous treatment effect analysis to understand when the LP-based model generates more revenue compared to the score-based model. Particularly, the results show that the LP-based algorithms generate more revenue when (multiple) purchases occur in an impression and when the recommended items have a high purchase-to-click ratio.
More specifically, for each impression $t$, we define $\textit{PurchaseNum}_{t}$ as the number of purchased items and $\textit{PurchaseToClick}_{t}$ as the ratio of the purchase number to the click number. Then we perform two OLS regressions with the following specifications:
$\textit{Revenue}_{t}$ $=\beta_{0}+\beta_{1}\textit{Approach}_{t}+\beta_{2}\textit{PurchaseNum}_{t}+\beta_{3}\textit{Approach}_{t}\times\textit{PurchaseNum}_{t}+\epsilon_{t},$ (14)
$\textit{Revenue}_{t}$ $=\beta_{0}+\beta_{1}\textit{Approach}_{t}+\beta_{2}\textit{PurchaseToClick}_{t}+\beta_{3}\textit{Approach}_{t}\times\textit{PurchaseToClick}_{t}+\epsilon_{t},$ (15)
where $\textit{Revenue}_{t}$ is the dependent variable; $\textit{Approach}_{t}$ is the binary variable indicating which method (score-based or LP-based) the current impression has been assigned to. In this analysis, we use LP90 for the LP-based method, because it obtains a more balanced improvement in terms of all three major metrics (revenue, GMV, Number of purchases).
Table 6: Heterogeneous Treatment Effect
Dependent varaible
Revenue Revenue
(1) (2)
LP90 0.004* -0.012***
[0.002] [0.004]
LP90 $\times$ PurchaseNum 0.304***
[0.115]
LP90 $\times$ PurchaseToClick 0.451****
[0.127]
Observations 1,381,222 1,381,222
  • Notes. The value inside the square brackets is the robust standard error.
    *$p<0.10$; ** $p<0.05$;***$p<0.01$;****$p<0.001$.
Table 6 presents the regression results from the heterogeneous treatment analyses. Column (1) shows that the coefficient of the interaction of LP90 approach and $\textit{PurchaseNum}_{t}$ is positive, meaning that the difference in financial performances between LP90 and Benchmark (the score-based algorithm) increases when there are more purchases occurred within a single impression. We believe this comes from the fact that the objective function in the LP problem $P_{1}$ maximizes expected revenue directly, in contrast to the Benchmark (score-based algorithm), which relies on a scoring function not directly tied to revenue. Further, Column (2) reports that the coefficient of the interaction of LP90 approach and $\textit{PurchaseToClick}_{t}$ is positive. Notice that an item with a high purchase-to-click ratio represents a high-quality item. It demonstrates that LP90 generates more revenue on high-quality items.
5.3 Revenue relevance trade-off
Through the LP model, we can identify the trade-off between revenue and relevance. When the parameter $\lambda$ in problem $P_{1}$ increases, the relevance increases as the optimal revenue decreases. In this section, we then take a closer look to inspect this trade-off with the empirical evidence from the field experiment. From September 13th, 2022 to September, 27th, 2022, we conducted a second A/B test on the same marketplace (dWeb), in which we mainly focused on exploring the comparison of performances for various values of $\lambda$. In total, we have set up four experiment groups, which are named respectively LP90, LP85, LP80, and LP75, taking the associated $\lambda$ to be 0.90, 0.85, 0.80, and 0.75. In total, we count 17,931,775 unique consumers and 199,681,466 impressions. Specifically, LP90 was assigned 4,483,091 consumers, generating 49,916,423 impressions; LP85 was assigned 4,487,531 consumers, generating 49,983,811 impressions; LP80 was assigned 4,479,883 consumers, generating 49,883,948 impressions; LP75 was assigned 4,481,270 consumers, generating 49,897,284 impressions. We concentrate on two metrics: RevenuePerConsumer and NumberPurchasesPerConsumer since these are the most essential evaluation metrics for revenue and relevance, as we explained before.
We set LP75 to be the benchmark, of which we compute the relative change of metrics for other groups over. Table 7 shows the apparent trade-off over these two matrics. For RevenuePerConsumer, compared to LP75, the relative changes of LP90, LP85, LP80 are 1.24%, 0.60%, and 0.30%. For the NumberPurchasesPerConsumer, compared to LP75, the relative changes of LP90, LP85, LP80 are -1.19%, -0.89%, and 0.11%. As $\lambda$, the hyperparameter controlling the overall relevance, decreases, NumberPurchasesPerConsumer drops as well while RevenuePerConsumer increases. Additionally, we plot the comparison of these metrics with the error bar in Figure 3.
(a) Relative change of number of purchases
(b) Relative change of revenue
Figure 3: Plots with the error bar showing the comparison of the relative change of (a) the NumberPurchasesPerConsumer and (b) RevenuePerConsumer for LP90, LP85, LP80, and LP75.
Table 7: Relative changes for the financial performance over LP75
LP90 LP85 LP80
RevenuePerConsumer (USD) (p-value) -1.19%  (0.12) -0.89%  (0.25) 0.11%  (0.89)
NumberPurchasesPerConsumer (p-value) 1.24%  (<0.01) 0.60%  (0.167) 0.50%  (0.25)
  • Notes. Most of the demonstrated p-values above do not display evidence for significant differences. The reason for that is largely attributed to the lack of data samples. The closer the mean of the two groups, the larger the amount of samples we would require to show the statistically significant differences. In practice, an A/B test would only run for at most two weeks, not being able to generate enough amount of samples. Nevertheless, instead of focusing on hypothesis testing, the purpose of this experiment is to explore the trade-off of the performances for different groups by changing the hyperparameter $\lambda$.
6 Extension: Overall planning
@@
6.1 LP-based algorithm for overall planning
@@
Step 1: Obtain dual variables for global constraints Given a historical dataset that records the impressions for each day, suppose there are $T$ impressions, $G$ unique consumers, and a group of $K$ selected sellers who pay additional Ad fees and accordingly pose requirements on inventory and revenue. For each impression $t$, it has $m_{t}$ slots and $n_{t}$ candidate items. Let $h_{it},v_{jt},r_{jt}$ denote the associated position weight for each slot $it$, revenue and relevance for each candidate item $jt$. Similar to the previous setting, for each impression $t$, we can get the maximum relevance score $a_{t}$ that any ranking can obtain, and define any relevance hyperparameter $\lambda_{t}\in[0,1]$.
@@
Suppose we have $T$ impressions in the historical data, indexed by $t$,i.e., $t=1,...,T$. We select $K$ unique sellers, indexed by $k=1,...,K$, and $G$ unique consumers, indexed by $g=1,...,G$. Each consumer $g$ has total $T_{g}$ impressions. Moreover, among these $T_{g}$ impressions, the candidate items are provided by $K_{g}$ sellers.
@@
6.2 Computational results for overall planning
In this section, we present the experiment results for the overall planning problem from Section 6. The overall planning is formalized as a large-scale problem using the data from a certain period; therefore, it can only be solved offline. Therefore, in this experiment, we formulate the optimization problem using the historical data of an entire day. Specifically speaking, we would set up the restrictions on the existing inventory $L$ and the targeted revenue $I$ for the selected top sellers who have the most listed items. In total, we filter out 5,408 sellers and formulate an optimization problem with 265,646,022 variables, 26,423,572 constraints, and 962,620,739 non-zeros. We use two solvers, which are Gurobi (version 9.0) and OR-Tools  [^3] https://developers.google.com/optimization/lp/lp_advanced, and compare their performances in terms of the solving time, number of iterations, and minimum memory spaces required.
@@ 6.2 Computational results for overall planning
Table 8: Experiment results for overall planning
@@
  • Notes. In total, the problem contains 265,646,022 variables, 26,423,572 constraints, and 962,620,739 non-zeros. We run each experiment with the same settings 5 times and take the average of each metric to report. In cases where the solving time limit is reached before the solution is given, we use the time limit(172,800 seconds) as its solving time in terms of calculating the average solving time.
  • *  For Primal and Dual simplex methods, Gurobi only allows 1 thread to solve the problem, which cannot be altered.
  • **  Dual simplex fails to give the correct solution, of which the metrics will not be reported.
  • ***  Only 1 out of 5 instances finished solving within the time limit.

### v3

@@ A.3 Proof of Theorem 1
    Stationarity: for all three subcases, we have $\boldsymbol{X}^{\text{LP}}\in\operatorname*{arg\,max}_{\boldsymbol{X}\in\mathcal{F}}L(\mu,\boldsymbol{X})$. In case B1 and B2, every split within the tied block maximizes $L(\mu^{*},\cdot)$, so our $\boldsymbol{X}^{\text{LP}}$ is one such maximizer; in case B3, since the order of $\boldsymbol{\delta}^{*}$ is strict over $j(1),\cdots,j(m)$, the stated permutation is the unique maximizer.
  4. 4.
    Complementary slackness: for all three cases, we have $\mu^{*}(\lambda a-\text{Rel}(\boldsymbol{X}^{\text{LP}}))=0$.
Since all KKT conditions are satisfied, therefore, $\boldsymbol{X}^{\text{LP}}$ is optimal for $P_{1}$. ∎
Then we have the following proof for Theorem˜1:
Proof.
Note that we could re-write $\mathbb{E}[\tilde{\boldsymbol{X}}]=(1-\alpha)\boldsymbol{X}_{+}+\alpha\boldsymbol{X}_{-}$. Then we discuss over each case.
Case A (Nonbinding relevance constraint). With $\mu^{*}=0$, we have $\boldsymbol{X}_{-}=\boldsymbol{X}_{+}=\boldsymbol{X}^{\text{LP}}$. Thus $\mathbb{E}[\tilde{\boldsymbol{X}}]=\boldsymbol{X}^{\text{LP}}$, where $\boldsymbol{X}^{\text{LP}}$ is the optimal solution to $P_{1}$.
Case B1 (Boundary tie between position $m$ and $m+1$). From Algorithm˜3, we see that both $\boldsymbol{X}_{-}$ and $\boldsymbol{X}_{+}$ have the exact same first rows m1m-1 with $\boldsymbol{X}_{+,i,j(i)}=\boldsymbol{X}_{-,i,j(i)}=1$ for all $i<m$. Thus, $\mathbb{E}[\tilde{\boldsymbol{X}}_{i,j(i)}]=\boldsymbol{X}^{\text{LP}}_{i,j(i)}=1$ for all $i<m$. In the $m$-th row, $\boldsymbol{X}_{+,m,j(m+1)}=1$ while $\boldsymbol{X}_{-,m,j(m)}=1$.
Thus, we get the $m$-th row of $\mathbb{E}[\tilde{\boldsymbol{X}}$] as:
$\mathbb{E}[\tilde{\boldsymbol{X}}_{m,j(m)}]=\alpha=\boldsymbol{X}^{\text{LP}}_{m,j(m)},$ (12)
and
$\mathbb{E}[\tilde{\boldsymbol{X}}_{mj(m+1)}]=1-\alpha=\boldsymbol{X}^{\text{LP}}_{m,j(m+1)}.$ (13)
Thus, $\mathbb{E}[\tilde{\boldsymbol{X}}]=\boldsymbol{X}^{\text{LP}}$.
B2 (Interior tie within the top $m$ positions). Similarly, we see that both $\boldsymbol{X}_{-}$ and $\boldsymbol{X}_{+}$ have the exact same rows except for $s$- and $s+1$-th row. Specifically, $\boldsymbol{X}_{-,s,j(s)}=1$ and $\boldsymbol{X}_{+,s+1,j(s+1)}=1$.
Therefore, we have $\mathbb{E}[\tilde{\boldsymbol{X}}_{s,j(s)}]=\mathbb{E}[\tilde{\boldsymbol{X}}_{s+1,j(s+1)}]=1-\alpha$ and $\mathbb{E}[\tilde{\boldsymbol{X}}_{s,j(s+1)}]=\mathbb{E}[\tilde{\boldsymbol{X}}_{s+1,j(s)}]=\alpha$. This gives: $\mathbb{E}[\tilde{\boldsymbol{X}}]=\boldsymbol{X}^{\text{LP}}$.
B3 (No tie within the top $m$ positions). Then the Lagrangian maximizer at $\mu^{*}$ is unique with $\boldsymbol{X}_{+}=\boldsymbol{X}_{-}=\boldsymbol{X}^{\text{LP}}$. Thus $\mathbb{E}[\tilde{\boldsymbol{X}}]=\boldsymbol{X}^{\text{LP}}$. ∎
A.4 Summary Statistics of the Sampled Data
Panel A in table A3 reports the summary statistics for our second dataset, introduced in Section 4.2, which consists of a random sample from the full-scale experiment. Panel B in table A3 presents the summary statistics for our third dataset, a random sample drawn prior to the launch of the field experiment. Both samples are at the impression level and we report summary statistics only for those variables whose disclosure is permitted by the company. All monetary variables are reported in USD.
Table A3: Summary Statistics.
Variable Mean Median SD Sample Size
Panel A: Random Sample of the Full-Scale Experiment
Average Product Price 2.69 0.31 57.33 2,071,833
Number of Clicks 1.07 1.00 0.53 2,071,833
Panel B: Random Sample Drawn Before the Experiment
Average Product Price 301.31 29.69 11223.03 13,370,159
Total Number of Candidate Items 70.17 60.00 47.90 266,003
Total Number of Items 11.34 12.00 2.35 266,003
Total Number of Sponsored Items 9.43 12.00 4.11 266,003
Number of Clicks 0.32 0.00 0.64 266,003
Seed Item Price 819.38 28.99 53,169.45 266,003
A.5 Extension: Overall Planning
@@
A.5.1 LP-Based Algorithm for Overall Planning
@@
Step 1: Obtain dual variables for global constraints. Given a historical dataset that records the impressions for each day, suppose there are $T$ impressions, $G$ unique consumers, and a group of $K$ selected sellers who pay additional Ad fees and accordingly pose requirements on inventory and revenue. For each impression $t$, it has $m_{t}$ slots and $n_{t}$ candidate items. Let $h_{it},v_{jt},r_{jt}$ denote the associated position weight for each slot $it$, revenue and relevance for each candidate item $jt$. Similar to the previous setting, for each impression $t$, we can get the maximum relevance score $a_{t}$ that any ranking can obtain, and define any relevance hyperparameter $\lambda_{t}\in[0,1]$.
@@
Suppose we have $T$ impressions in the historical data, indexed by $t$,i.e., t=1,,Tt=1,...,T. We select $K$ unique sellers, indexed by k=1,,Kk=1,...,K, and $G$ unique consumers, indexed by g=1,,Gg=1,...,G. Each consumer $g$ has total $T_{g}$ impressions. Moreover, among these $T_{g}$ impressions, the candidate items are provided by $K_{g}$ sellers.
@@
A.5.2 Computational results for overall planning
In this section, we present the experiment results for the overall planning problem from Section A.5. The overall planning is formalized as a large-scale problem using the data from a certain period; therefore, it can only be solved offline. Therefore, in this experiment, we formulate the optimization problem using the historical data of an entire day. Specifically speaking, we would set up the restrictions on the existing inventory $L$ and the targeted revenue $I$ for the selected top sellers who have the most listed items. In total, we filter out 5,408 sellers and formulate an optimization problem with 265,646,022 variables, 26,423,572 constraints, and 962,620,739 non-zeros. We use two solvers, which are Gurobi (version 9.0) and OR-Tools  [^1] https://developers.google.com/optimization/lp/lp_advanced, and compare their performances in terms of the solving time, number of iterations, and minimum memory spaces required.
@@ A.5.2 Computational results for overall planning
Table A4: Experiment Results for Overall Planning
@@
  • Notes. In total, the problem contains 265,646,022 variables, 26,423,572 constraints, and 962,620,739 non-zeros. We run each experiment with the same settings 5 times and take the average of each metric to report. In cases where the solving time limit is reached before the solution is given, we use the time limit(172,800 seconds) as its solving time in terms of calculating the average solving time. *   For Primal and Dual simplex methods, Gurobi only allows 1 thread to solve the problem, which cannot be altered. **   Dual simplex fails to give the correct solution, of which the metrics will not be reported. ***   Only 1 out of 5 instances finished solving within the time limit.